OpenStudy (anonymous):

Can somebody factor the following? 1) x^3 - x^2 + x - 2 = 0 2) x^2 + x -1 = 0

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

x^3-x^2+x-2=x^2+x-1 x^3-x^2-X^2+x-x-2+1=0 x^3-2x^2-1=0 x1=2.20556, x2=-0.10278

OpenStudy (anonymous):

and do you know how to factorize (break it down) : x^3 + 10 ? in this type of answer (x +..) and ( x-..) ?